OpenStudy (anonymous):

medals rewarded! Solve the matrix equation. X+[-16] [ -9] + 16 = -29 [ -12] 9 -29 [ 7 ] -6 -10 1 26

OpenStudy (anonymous):

do you mean ... \[X+\left[\begin{array}{c} -16\\ -9\\ -12\\ 7 \end{array}\right]+\left[\begin{array}{c} 16\\ 9\\ -6\\ 1 \end{array}\right]=\left[\begin{array}{c} -29\\ -29\\ -10\\ 26 \end{array}\right]\]

OpenStudy (anonymous):

yeah, sorry it didn't come out right

OpenStudy (anonymous):

\[X=-\left[\begin{array}{c} -16\\ -9\\ -12\\ 7 \end{array}\right]-\left[\begin{array}{c} 16\\ 9\\ -6\\ 1 \end{array}\right]+\left[\begin{array}{c} -29\\ -29\\ -10\\ 26 \end{array}\right]\] now just add

OpenStudy (anonymous):

what would that come out to

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

it can't be as simple as adding, -16 + 16 = 0. then what?

OpenStudy (zarkon):

-(-16)-16+-29

OpenStudy (zarkon):

that would be the first entry of \(X\)

OpenStudy (anonymous):

so the second would be -(-9)-9+-29

OpenStudy (anonymous):

@Zarkon

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

that would come out to -29 and that can't be right.

OpenStudy (zarkon):

if that is not correct then make sure you typed the original question correctly